### v1

# Osmolyte-Modulated Differential Capacitance and Disjoining Pressure for Nanoconfined Electrolytes: A Modified Poisson-Boltzmann Theory

Abstract
This study employs modified Poisson-Boltzmann theory to systematically investigate the influence of zwitterionic osmolyte additives to an electrolyte solution on disjoining pressure and electric differential capacitance within charged slit-like nanopores with conductive walls. We demonstrate that increasing concentrations of zwitterionic osmolytes result in a marked synergistic enhancement of both disjoining pressure and differential capacitance, highlighting their dual role in improving supercapacitor performance. The insights gained underscore the unique capabilities of zwitterionic osmolytes as multifunctional additives for fine-tuning the properties of electric double layers, thereby bridging the gap between capacitive efficiency and electrode longevity.
I Introduction
Liquid-phase electrolytes confined within nanoscale porous electrodes are central to advancing electrochemical energy storage technologies, particularly supercapacitors [1, 2, 3, 4]. These systems rely on the electric double layer (EDL) formed at the interface between a conductive porous electrode and an electrolyte, where rapid ion adsorption enables high power density and cyclability. In supercapacitors, the EDL’s structure within carbon micropores [3, 5, 6] dictates critical performance metrics such as energy storage capacity and differential capacitance – a key parameter defining the charge stored per unit voltage [7, 8]. The nanoconfinement of ions in these pores amplifies interfacial phenomena, where steric effects, ion correlations, and dielectric saturation diverge sharply from bulk behavior, directly influencing the EDL’s thermodynamics and, consequently, the device’s efficiency [9, 10].
A paramount challenge in optimizing supercapacitors lies in quantifying the disjoining (or, more generally, solvation [11]) pressure (stress) – the mechanical force per unit area exerted by confined electrolytes on pore walls [12]. This pressure governs the structural stability of porous electrodes during charge-discharge cycles. Cyclical swelling pressures from ion intercalation can induce electrode deformation, pore collapse, or fracture, leading to capacity fade and device degradation [13, 14, 15, 16, 17]. Moreover, the disjoining pressure modulates ion packing density within pores [9, 18, 19], directly impacting the EDL’s differential capacitance.
Accurately predicting disjoining pressure remains complex due to competing nanoscale effects, including electrostatic correlations, solvation forces, and ion size limitations, which traditional mean-field models like Poisson-Boltzmann (PB) theory fail to capture. Standard PB frameworks, while foundational for modeling EDLs, neglect finite ion size [7, 20, 21, 22], dielectric saturation [23, 24, 25, 26, 27, 28], ion solvation effects [29, 30, 28], electrostatic correlations [31, 32, 33, 34, 35, 36, 37], structural effects [38, 39, 40], short-range specific interactions of species [41, 27, 28], and internal electric structure of ions and solvent molecules [42, 43, 44], and, finally, specific adsorption of ions [45, 32] rendering them rather inaccurate for nanoconfined systems. Modified Poisson-Boltzmann (MPB) equations address these gaps by incorporating non-mean-field corrections, asymmetric ion partitioning, and surface interactions (see, [10, 46, 47] and references therein). For supercapacitors, these models based on the MPB theory may be essential for designing porous electrodes with optimized pore geometries and surface chemistry to maximize the differential capacitance [9, 48].
A critical yet underexplored issue is the role of zwitterionic additives – molecules bearing spatially separated positive and negative charges – in modulating disjoining pressure and EDL thermodynamics and dynamics. In energy storage contexts, zwitterions could enhance ion accessibility within micropores by preventing aggregation and optimizing charge distribution, thereby increasing differential capacitance. Their dipolar nature introduces localized charge regulation [49, 50] and screening effects, potentially reshaping confinement thermodynamics. While zwitterions are established as biocompatible stabilizers in biological systems (e.g., osmolyte-mediated membrane stabilization [51, 52]), their impact on electrochemical systems remains speculative. Specifically, their ability to alter disjoining pressure and ion structuring in nanopores could unlock strategies to mitigate electrode degradation and enhance differential capacitance.
In a recent paper [53], the authors examined the theory of EDL that forms at the metal-solution interface of zwitterionic polymers. These unique polymers feature oppositely charged sites linked by flexible, neutral linear chains. By employing self-consistent field theory, the authors revealed that, from the perspective of supercapacitor applications, these systems concurrently exhibit the properties of electrolytes and dielectric materials. This dual nature leads to the formation of EDL and facilitates the passage of an electric field through the system. Furthermore, the authors speculated on the potential of hybrid supercapacitors that incorporate salt solutions along with zwitterionic polymer additives. Such devices could function effectively at low electrode potentials, where ion movement significantly contributes to capacitance formation, and at high potentials, where the movement of zwitterions enhances capacitance. Additionally, it is important to consider low-molecular-weight zwitterionic co-solvents, particularly betaine oligopeptides. This is crucial because these relatively small molecules can more easily penetrate the nanoscale pores of supercapacitor electrodes than polymeric ones with high conformational entropy [54]. In addition, due to their significant dipole moments, betaine molecules could sharply increase the differential capacitance when they are present in electrolyte solutions, due to the enhancement of the local dielectric permittivity [55, 56, 57, 27, 58, 59, 10]. Consequently, investigating electrolyte solutions that contain admixtures of these zwitterionic molecules, especially in the confined geometries of electrified nanopores, is an essential task for advancing the fields of modern chemical physics and electrochemistry.
This study uses the MPB formalism to investigate the disjoining pressure and the differential capacitance behavior of the EDL in charged slit-like nanopores filled with electrolyte solutions containing additives of a zwitterionic co-solvent. We will show that increasing the concentration of zwitterions increases the electrostatic disjoining pressure. Importantly, we will demonstrate that there is a simultaneous increase in the differential capacitance when zwitterions are added to an electrolyte solution. This has important implications for the design of supercapacitors.
II Theoretical background
Let us consider a 1:1 electrolyte solution with an admixture of the zwitterionic molecules confined within an infinite slit-like pore, with electrified conductive walls that are separated by a distance of a porous electrode (Fig. 1). We assume that the confined solution is in equilibrium with the bulk solution, so that the grand thermodynamic potential per unit wall’s surface area can be written in the form [47]
where the first term in the right hand side of eq. (1) is the electrostatic free energy of pure electrolyte solution (without zwitterions) in the mean-field approximation with being the local electric field; are the local ionic concentrations; is the local electrostatic potential and is the local charge density of ions with charges ; is the dielectric constant of the solvent which we model as continuous dielectric medium and is the permittivity of vacuum. The second term is the free energy of dipole gas in "external" self-consistent field [27, 58, 59]; is the local concentration of zwitterions with dipole moment . The third term is the grand thermodynamic potential of solution within the reference fluid [47] model taking into account the interaction of the species with the walls; is the Boltzmann constant, is the temperature. In this paper we use the multicomponent asymmetric ideal (without attraction) lattice gas model with the following free energy density [21, 38, 40, 60]
where , is the volume fraction of species (), is the cell volume of the cubic lattice; represents the number of lattice cells occupied by species .
The self-consistent field equations can be obtained from the grand thermodynamic potential variation, i.e. from the following Euler-Lagrange equations
which yield the Poisson equation
with the local permittivity of the solution [56, 61]
where is the Langevin function, and the chemical equilibrium equations for the local concentrations of species
where
is the intrinsic chemical potential of the species within the asymmetric lattice gas model and is the total potential energy of interaction between species and the walls; in turn is the single-wall potential which will be specified below. The bulk chemical potentials, , are determined by eqs. (8), with and , where is the ion concentration in the bulk and is the bulk zwitterion concentration.
To obtain the local concentrations of ions and zwitterions, we need to solve a system of coupled equations (4) and (6), 7), with the boundary conditions , where is the surface potential.
To calculate the differential capacitance, we need to know the wall’s surface charge density, ,where is the permittivity on the wall surface. The differential capacitance is .
The disjoining pressure can be calculated by [61, 62]
where is the normal pressure at generated by the species. The local osmotic pressure is and is the bulk pressure, i.e. pressure, , at . For a relatively large pore, where the single-wall potential range is much smaller than the pore width, as is the case in this study, we can ignore the integral on the right-hand side of eq. (9) and use the Langmuir equation [12, 61], , to calculate the disjoining pressure.
III Numerical results and discussions
In this section, we present the results of our calculations regarding the electrical differential capacitance and disjoining pressure for a confined aqueous () electrolyte containing zwitterionic additives. These results are derived from the self-consistent field equations (4), (6), and (7). We will focus on zwitterionic glycine in both peptide and dipeptide forms. For glycine as a peptide, we assume that each molecule occupies a single cell () and has a dipole moment of (for details, see [63]). In the case of dipeptide glycine, we assume while assuming the dipole moment to be (see also [63]). Additionally, we assume that both cations and anions occupy only a cell, i.e., . We will also consider the case of specific adsorption of ions modelling the single-wall potentials as the Gaussian-well potentials [40]
where is the energetic parameter of interaction of species with the electrode and is the corresponding interaction range.
Note that in this study we consider sufficiently wide pores ( nm), as in this case, we may not need to take into account the contribution of solvation forces to the disjoining pressure associated with the strong heterogeneity of the solvent (water) [11, 64, 65]. Additionally, we consider dilute solutions of electrolytes and zwitterions, so that non-local effects that are important for ionic liquids and concentrated solutions [39, 38, 40], can be ignored.
III.1 Differential capacitance
Fig. 2 shows the differential capacitance profile of the confined aqueous electrolyte in a pore with a thickness of nm for a rather dilute electrolyte concentration M, at different glycine concentrations in both peptide and dipeptide forms. As can be seen, an increase in the concentration of peptide and dipeptide glycines leads to a slight increase in differential capacitance at relatively low surface potentials ( V) and a more significant increase at higher concentrations ( V). Note that the effect persists for moderately concentrated electrolyte solutions (see Fig. 3). Note that we only show data for one pore size, nm, as the differential capacitance changes only slightly with pore size at nm. Furthermore, changes in pore size in this range do not violate the effect of differential capacitance enhancement discussed.
The pronounced increase in differential capacitance observed in the high-potential region can be attributed to several key factors. Near the electrode interface, the charge from the electrode is effectively screened by mobile ions, resulting in a highly inhomogeneous electric field that exerts a significant influence on polar glycine molecules. This electric field induces a dielectrophoretic force that effectively draws the glycine molecules toward the electrode surface [57, 27, 59]. As a consequence, there is an enhanced concentration of glycine near the electrode (as illustrated in Fig. 4), which subsequently increases the local permittivity of solution. This increase in permittivity contributes to a greater accumulation of ionic charge at elevated surface potentials within the EDL, driven by the electrostatic attraction between ion charge and dipole moments of the glycine molecules (see Fig. 1). Consequently, this phenomenon results in a larger differential capacitance.
Indeed, as demonstrated in Figure 5, the concentration of cations at the electrode monotonically increases with the bulk peptide glycine concentration. This effect is relatively negligible at low surface potentials, but becomes pronounced when the potential exceeds a certain threshold (approximately ( V). It is important to note that a similar enhancement occurs when considering the specific adsorption of cations onto the electrode surface (refer to Fig. 6). Our calculations suggest that the charge enhancement in the EDL, induced by the additive of zwitterionic osmolytes into the electrolyte solution, is a generic physical phenomenon that can be qualitatively interpreted within the framework of the MPB theory.
It is important to mention that this study yields non-physically large values for the differential capacitance. This discrepancy arises because our theoretical framework does not incorporate important physical factors, such as the explicit molecular structure of the solvent, specific adsorption of water molecules onto the electrode, and ion solvation effects, and ion-solvent specific interactions [28]. One more reason consists in rather rough accounting of the excluded volume effect of the species via the asymmetric lattice gas model. In this paper, our primary objective is to illustrate the influence of introducing zwitterions into a solution through a simplified model, without prioritizing alignment with experimental data.
III.2 Disjoining pressure
Recent experimental studies by Govrin et al. [49] reveal that incorporating osmolytes into an electrolyte solution diminishes the electrostatic screening of charged surfaces by the mobile ions present in the solution, thereby amplifying the repulsive forces between them. In the context of slit-like pores with charged walls, this phenomenon is expected to result in an enhancement of the electrostatic component of the disjoining pressure as the concentration of zwitterionic molecules in the bulk solution increases. The MPB theory predicts this behavior, as illustrated in Figures 7 and 8. It suggests that the disjoining pressure will rise with increasing glycine concentration, applicable to both its peptide and dipeptide forms. Notably, this effect is observed at two distinct ranges of surface potentials: at low surface potentials, where the glycine concentration within the pore resembles that of the bulk solution, and at higher potentials, where the glycine concentration near the wall exceeds that of the bulk. In both scenarios, the impact is more pronounced for glycine dipeptides compared to peptides, as depicted in Figures 7 and 8. This enhanced effect can be attributed to a substantial increase in the local permittivity of the solution, which reduces walls’ charge screening by the mobile ions and intensifies electrostatic repulsion of the charged walls. In the case of relatively large surface potentials, where an increase in dipeptide glycine concentration leads to a significant increase in differential capacitance, as well as weakening of electrostatic screening in pores ( nm), the disjoining pressure increases due to the fact that the local glycine concentration within these pores is higher than in the surrounding solution. This leads to a higher osmotic pressure at the center of the pore compared to that in the bulk solution, not only due to the ionic contribution, as in the case of small potentials, but also due to the presence of excess osmolyte molecules.
IV Discussion
Now, we would like to discuss possible implications of the predicted effects for practical applications.
First, it is particularly noteworthy that incorporating zwitterionic osmolytes into an electrolyte solution can lead to a substantial enhancement in differential capacitance, even at low electrolyte concentrations. This characteristic presents a significant advantage in electrochemical applications, as it effectively reduces unwanted electrochemical activity of ions on the electrode surface, without loss of differential capacitance. By partially substituting these ions with zwitterionic osmolytes the overall electrochemical performance of the system is improved. Their neutral charge prevents participation in Faradaic reactions, thereby reducing parasitic electrochemical activity (e.g., electrode corrosion) that typically plagues traditional electrolytes. This reduction in ionic activity could be beneficial as it minimizes adverse side reactions that can negatively affect the longevity and stability of the electrodes.
Second, during the operation of the supercapacitors, the micropores found within the electrodes are susceptible to collapse due to van der Waals forces acting between the pore walls, as well as chemical interactions with the ions present in the solution. Such structural degradation can severely compromise the performance and efficiency of the supercapacitor. In contrast, the osmolytic effects brought about by the zwitterionic osmolyte can enhance the repulsive forces acting between the pore walls, thereby reducing the likelihood of collapse and preserving the structural integrity of the micropores. Maintaining this integrity may be crucial, as it ensures that the porous electrodes retain a high surface area available for charge storage. A higher surface area correlates directly to improved capacitance and energy storage capabilities. Consequently, the introduction of zwitterionic osmolytes not only stabilizes the pore structure but also contributes to maintaining consistent capacitance levels over extended periods, ultimately enhancing the reliability and performance of supercapacitors in various applications. Thus, the strategic use of zwitterionic osmolytes can represent a significant advancement in electrolyte formulations, leading to more efficient energy storage devices.
V Conclusion
In conclusion, we utilized modified Poisson-Boltzmann theory to systematically explore the role of zwitterionic osmolyte additives in modulating the disjoining pressure and differential capacitance of electric double layers within charged slit-like nanopores with conductive walls. Our findings revealed that increasing zwitterion concentration induces a pronounced synergistic enhancement of both disjoining pressure and differential capacitance, a dual effect critical for optimizing supercapacitor performance. We believe that the obtained insights may indicate on the unique potential of zwitterionic osmolytes as multifunctional additives for tailoring electric double layer properties, bridging the gap between capacitive efficiency and electrode durability. By elucidating the molecular mechanisms underlying this synergy, our work provides a foundational framework for designing advanced supercapacitors that harmonize high energy density with long-term operational stability. Future research could expand on these insights by exploring zwitterion-specific interactions in diverse electrolyte systems and nanopore geometries, further discovering their transformative potential for modern energy storage technologies.
Data availability statement. The data that supports the findings of this study are available within the article.
Acknowledgments
The authors thank the Russian Science Foundation (Grant No. 24-11-00096) for financial support.